4.18 A random sample of size 25 is selected from a population with mean μ = 85 and standard deviation σ-4. Approximate the following probabilities using the central limit theorem (a) PrX 86, 6451 (b) PrX < 84.340] (c) Pr(83.04 〈 X < 86.96]
Answer #1

P(small ar{X} < A) = P(Z < (A - small mu_{ar{x}} )/small sigma_{ar{x}})

Here, small mu = 85

small sigma = 4

n = 25

By central limit theorem, small mu_{ar{x}} = small mu = 85

small sigma_{ar{x}} = small sigma/sqrt{n}

= 4/V25

= 0.8

(a) P(small ar{X} > 86.645) = 1 - P(small ar{X} < 86.645)

= 1 - P(Z < (86.645 - 85)/0.8)

= 1 - P(Z < 2.06)

= 1 - 0.9803

= 0.0197

(b) P(small ar{X} < 84.340) = P(Z < (84.340 - 85)/0.8)

= P(Z < -0.825)

= 0.2047

(c) P(83.04 < small ar{X} < 86.96) = P(Z < 86.96) - P(Z < 83.04)

= P(Z < (86.96 - 85)/0.8) - P(Z < (83.04 - 85)/0.8)

= P(Z < 2.45) - P(Z < -2.45)

= 0.9929 - 0.0071

= 0.9858